### v1

@@ 2.2 Instruments used for the observations and data reduction
  • UVES (Ultraviolet and Visual Echelle Spectrograph [^6] https://www.eso.org/sci/facilities/paranal/instruments/uves.html), attached at the UT2 (Unit Telescope 2) of VLT (Very Large Telescope), placed at Paranal (Chile). The red arm was used, equipped with the grism CD#3, covering the wavelength interval 4760–6840 Å, and the central wavelength at 5800 Å. The 1 arcsec slit, which provides a dispersion of Rsimilar-$\sim$47,000, was selected for all the targets.
  • HARPS-N (High Accuracy Radial velocity Planet Searcher in North hemisphere [^7] https://www.tng.iac.es/instruments/harps/), attached at the 3.5m Telescopio Nazionale Galileo (TNG). HARPS-N features an echelle spectrograph covering the wavelength range between 3830 to 6930 Å, with a spectral resolution R=115,000.
  • ESPaDOnS (Echelle SpectroPolarimetric Device for the Observation of Stars [^8] https://www.cfht.hawaii.edu/Instruments/Spectroscopy/Espadons/) attached at the 3.6m Canada-France-Hawaii Telescope (CFHT). ESPaDOnS provides a spectral resolution R=81,000 in the wavelength range between 3700 to 10500 Å.
Typically we obtained 1-2 epoch spectra with UVES, 2-3 with CFHT and three or more with HARP-N. The Signal-to-Noise ratio (SNR) was larger than 40 for about 80% of the spectra. In total, in this work, we analysed 331 spectra for 180 DCEPs. A complete list of the time of acquisition and the individual SNR values is reported in Table 2.
Reduction of all the spectra, which included bias subtraction, spectrum extraction, flat fielding and wavelength calibration, was done by automatic pipelines provided by the three instrument teams so that we downloaded the science-ready 1-dimensional spectra (for more details on the data reduction of HARPS-N, UVES and ESPaDOnS spectra, see Ripepi et al., 2021; Trentin et al., 2023; Bhardwaj et al., 2023, respectively).
Finally, from Table 1 it can be noted that the stars OGLE-GD-CEP-0026, OGLE-GD-CEP-1278 and OGLE-GD-CEP-1290 have been observed with both the UVES and HARP-N instruments. These stars are therefore useful for cross-checking of the results obtained with the different instruments (see Sect. 4).
Figure 2: Example of spectral synthesis for our targets in the region from $\lambda$  = 6110 Å to $\lambda$  = 6151 Å. The colours differ according to the particular instrumental equipment, specifically in red the UVES spectra, in green the HARPS spectra, and in blue the ESPADONS spectra. The spectra are ordered, from top to bottom, by decreasing metallicity. The main spectral lines have been identified at the bottom.
@@
Figure 4: Histograms of the distribution of the chemical abundances derived in this study. Gaussian fits have been over imposed (red dashed lines) with the respective mean value and FWHM reported in each panel.
Figure 5: Chemical elements in the form of [X/Fe] plotted against the Iron content. The colors are the same as Fig.2, with the addition of the targets from R21 and T23 in orange and yellow, respectively (see also Fig.6 and 9 for the color legend).
Figure 6: $\alpha$-elements abundance in the form of [$\alpha$/Fe] plotted against the Iron content. The colour pattern, the same as Fig. 5, is described in the upper right legend. The dotted blue horizontal lines highlight [$\alpha$/Fe] abundances at 0.4, 0.2 and 0 dex.
@@ 3 Spectroscopy
A widely adopted tool in the literature for estimating effective temperature is the line depth ratios (LDR) method (Kovtyukh & Gorlova, 2000). This method offers the advantage of being sensitive to temperature variations while remaining unaffected by abundance changes and interstellar reddening. Typically, for each spectrum in our targets, we measured approximately 32 LDRs.
For the remaining parameters ($\xi$ and log g), we employed an iterative approach. The estimation of microturbulence involved ensuring that iron abundances showed no dependence on equivalent widths (EWs), i.e., that the slope of [Fe/H] against EWs is null. To achieve this occurrence, we initially measured the equivalent widths of 145 Fe I lines using a custom Python semi-automatic routine. The line sample was extracted from the list published by Romaniello et al. (2008), and the routine minimized errors in continuum estimation on the spectral line wings. EWs were converted to abundances using the WIDTH9 code (Kurucz & Avrett, 1981), applied to the corresponding atmospheric model calculated using ATLAS9 (Kurucz, 1993). In this calculation, the influence of $\log$ g was not considered, as neutral iron lines are insensitive to it. Subsequently, surface gravities were estimated through a similar iterative procedure, imposing ionization equilibrium between Fe I and Fe II. The adopted list of 24 Fe II lines was extracted from Romaniello et al. (2008). To circumvent issues arising from spectral line blending due to line broadening, a spectral synthesis technique was applied to our spectra. Synthetic spectra were generated in three steps: i) plane-parallel local thermodynamic equilibrium (LTE) atmosphere models were computed using the ATLAS9 code (Kurucz, 1993), employing the stellar parameters; ii) stellar spectra were synthesized using SYNTHE (Kurucz & Avrett, 1981); iii) the synthetic spectra were convoluted to account for instrumental and line broadening. This convolution was evaluated by matching the synthetic line profiles to a selected set of observed metal lines.
For a total of 29 different chemical elements, it was possible to detect the spectral lines used for the estimation of the abundances. For all elements we performed the following analysis: we divided the observed spectra into intervals, 25 Å or 50 Å wide, and derived the abundances in each interval by performing a $\chi^{2}$ minimization of the differences between the observed and synthetic spectra. The minimization algorithm was written in Python, using the amoeba routine.
@@
The adopted lists of spectral lines and atomic parameters are taken from Castelli & Hubrig (2004), who have updated the original parameters of Kurucz (1995). When necessary, we also checked the NIST database (Ralchenko & Reader, 2019). In Fig.2, we plot an example of some of the spectra analysed in this work.
More details about the spectral lines that can be typically detected and used to estimate chemical abundances can be found in R21 and T23.
The final list of abundances for all the stars is reported in Table 3.
Figure 7: Example of 2 spectra around the Sc II line at $\lambda\lambda$ $\AA$ for the stars BQ Vel (in yellow from the T23) and ATLAS-J104.8954-08.3771 (in red from the UVES sample analysed in this paper). Other spectral lines are highlighted with dashed black lines. Atmospheric parameters and metallicity differences between the two targets are reported.
Figure 8: Example of 2 spectra around the Zn I line at $\lambda\lambda$ $\AA$ for the stars OGLE-GD-CEP-0228 (in yellow from the T23) and OGLE-GD-CEP-0048 (in red from the UVES sample analysed in this paper). Other spectral lines are highlighted with dashed black lines. Atmospheric parameters and metallicity differences between the two targets are reported.
@@ 4 Results
As a first step into the analysis of our results, we verified that for the three targets observed with both UVES and HARPS-N and mentioned in Sect. 2.2 we find consistent results among the estimated abundances, ruling out the possibility of a bias among spectra from different instruments. As showed in Fig.3, for all three stars we find comparable chemical values between the UVES and HARPS spectra and no systematics are detected.
@@
In Fig.4 we show histogram distributions of each derived element abundance, for a total of 29 different chemical species. Bins have been fixed to 0.15 dex reflecting typical error values. Each distribution has been fitted with a Gaussian curve. It is worth noting that most of these Gaussian curves have FWHM ¿ 0.25 dex, which means that for these elements we have an almost homogeneous distribution over a broad range of values. Moreover, when confronting current Fig.4 with Fig.3 of T23, it is possible to appreciate how this new sample better covers the whole range of abundances without sharp ”jumps” among bins that appeared in T23 for elements (see for example the O, Ca, and Fe panels). We complemented our sample with those presented in the previous C-MetaLL papers R21 and T23 reaching a total of 292 pulsators analysed homogeneously.
In Fig. 5 we plot each element (in its [X/Fe] form) in relation to iron. Generally, we confirm all the trends already found in R21 and T23 where we divided the chemical species into groups according to their main channel of formation within the stellar evolution models. In more detail:
@@
  • For light and $\alpha$ elements (from C to Ti) we observe a descending behaviour at lower metallicities until [Fe/H] $\approx$ -0.5 dex, with a flattening effect at higher values. Moreover, we note that while in T23 the region with over solar abundances at lower [Fe/H] was poorly populated with stars that appeared as outliers (yellow points), in this work we can confirm a negative trend that flattens at higher iron abundances. In Fig. 6 we plotted the $\alpha$ elements vs. [Fe/H], confirming the $\alpha$ enhancement at lower metallicities that had already been observed in literature from both observational (Hayden et al., 2015; Duffau et al., 2017; Trentin et al., 2023) and theoretical (Tinsley, 1979; Matteucci & Brocato, 1990; Palicio et al., 2023) context. We note that for the Sc element, most of the low metallicity stars studied in this paper appear systematically less abundant (around solar values) and with a flatter distribution than those presented in T23 which are clearly over-abundant and with a slight descending trend towards higher [Fe/H] values. In order to investigate this discrepancy between the two samples we selected stars with similar atmospheric parameters and (and possibly similar [Fe/H] abundances) and compared the spectra around the available scandium lines. An example with the two targets BQ_Vel from T23 and ATLAS-J104.8954-08.3771 is shown in Fig. 7. It was indeed found that stars from T23 show more intense Sc lines (that is, higher abundances) than those presented in this work. This suggests that this effect could be real and not caused by systematics in the analysis. Possible effects due to different instruments are also excluded since most of the stars involved were observed with UVES.
@@
  • Iron peak elements (from V to Zn) present a descending trend over the range of metallicities, with the exceptions of Mn and Zn which appear flatter. It is worth noting how for the latter element, the objects presented in this work, observed with UVES (red points), populate the over-solar and solar abundances zone at lower metallicities, while most of the previous targets from T23 (also observed with UVES) were found to be under-solar. This is a similar but opposite behaviour to the Sc case. In the wavelength range covered with the UVES instrument, however, there’s only one available neutral zinc line at $\lambda\lambda$ 6362.338 Å. Although this line is weak and it could be difficult to detect when the S/N is not optimal, a difference in the line depth has indeed been observed as in the case of Sc (see Fig. 8). A similar study has been performed for the Copper, for which 2 neutral lines at $\lambda\lambda$ 5105.5 and 5218.201 Å. In this case the UVES targets are slightly under-abundant with respect to the other samples, but again a visual inspection confirmed this difference.
@@
  • For heavier elements (From Sr to Gd) we either find a descending behaviour or a sort of arch-form trend with a bump at [Fe/H] $\approx$--0.2 dex, after which we find that abundances decrease with increasing metallicities. This curved trend was already found in the literature related to Open Clusters studies (e.g. Molero et al., 2023) and is typical of those elements produced via the s-process. While the descending branch at higher metallicities is caused by the iron production from SNeIa, at lower metallicities pollution from low and intermediate-mass stars during the AGB phase cause the rising trend. Given the low number of stars, we cannot draw a firm conclusion for both Sr and Sm. As happened for Sc, Cu and Zn, also for the zirconium (Zr II line at $\lambda\lambda$ 6114.853Å ) there seems to be a sort of bifurcation at the low-metallicity tail.
@@
Figure 9: Galactic iron radial gradient. The colours are the same as Fig. 5. The black line represents the linear fit computed. Results of the fit are reported both in the figure and in Table 9.
@@
Figure 10: Galactic radial gradient for all the estimated chemical elements. The lines and colors are the same as Fig. 9. The linear coefficients are listed in Table 9
@@ 4.1 Radial gradients and spatial distribution
Following the procedure described in T23 we estimated the Galactocentric distances $R_{GC}$ and estimated the radial gradients for all the available chemical species. All fits have been carried out with the python LtsFit package (Cappellari et al., 2013), which allows one to use weights on both axes and implement a robust outlier removal. In Fig. 9 we plotted the iron radial Galactic gradient. As expected, we found a clear negative trend. Moreover, an almost homogeneous distribution in metallicity is reflected in a homogeneous distribution in terms of distance, with the stars spanning a Galactocentric distance between 5 and 20 kpc. One first difference with our previous work is the lack of a visible break at around 9.25 kpc. This is confirmed when trying to compute a two-line fit which overlaps almost perfectly with the single-line case. Also from the quantitative point of view, we noted a significant discrepancy in the value of the slope. In more detail, we found a higher value (in an absolute sense) compared with the single-line case and in between the inner and outer slope of the two-line fit. On the other side, when considering the results in T23 obtained after a binning division of the sample, to compensate for the non-uniform distribution, the overall slope is in agreement with our most recent study (there is a good agreement with the two-line fit as well). This puts a further emphasis on the importance of having access to a sample of stars as uniform and homogeneous as possible in terms of distance (and metallicity), one of the key points of the C-MetaLL project. In Fig. 10 we show the radiant galactic gradient for the other 28 elements. The estimated coefficients together with the rms, are listed in Table  9. These results generally agree with those listed in T23. As in our previous work, Ba has the largest dispersion. Other elements (i.e. Sc, Ti and Zn) present instead a difference $>$0.01 dex kpc-1. For these cases we tried to repeat the linear regression including also the literature sample used in T23, obtaining results in perfect agreement with those of this work, once again highlighting the role of the farthest (and most metal-poor) objects and their weight in the linear fit. This is further remarked when considering the significant difference in the number of objects used in the two works, counterbalanced by the homogeneity of the analysis and the sampling of the stars along the Galactocentric radii.
@@
Taking into consideration this information, we can find a possible justification for the anomalous behaviour of the DCEPs studied here concerning the Sc, Cu, Zn and Zr abundances which have been discussed in the previous paragraph. As showed in Fig. 12, the evident and abrupt abundance change for the outermost stars could be because these two ”blocks” of stars are tracing two different spiral arms (Outer and OSC) and not only the Outer arm as stated in our previous work. No firm conclusion could be drawn about the opposite behaviour of Zn compared with the other 3 elements.
Figure 11: Galactic polar distribution of the stars. The new targets presented in this work are highlighted with black borders. The points are colour-coded according to [Fe/H] values and specified on the colour bar. The position of the sun is shown with a yellow-black symbol. Spiral arms from Reid et al. (2019) are over imposed. The colours are explained in the upper-right legend.
Figure 12: Galactic polar distribution of the stars. The new targets presented in this work are highlighted with black borders. The points are colour-coded according to the chemical element specified on the colour bar. The position of the sun is shown with a yellow-black symbol. We warn the reader about the different colour bar scales for each element.

### v3

@@ 2.2 Instruments used for the observations and data reduction
  • The Ultraviolet and Visual Echelle Spectrograph (UVES, Dekker et al., 2000)666 https://www.eso.org/sci/facilities/paranal/instruments/uves.html), attached at Unit Telescope 2 (UT2) of Very Large Telescope (VLT), placed at Paranal (Chile). The red arm was used, equipped with the grism CD#3, covering the wavelength interval 4760–6840 Å, and with the central wavelength at 5800 Å. The 1 arcsec slit, which provides a dispersion of Rsimilar-$\sim$47,000 (sampling 2 px), was selected for all the targets.
  • The High Accuracy Radial velocity Planet Searcher for the Northern hemisphere (HARPS-N, Mayor et al., 2003; Cosentino et al., 2012)  [^7] https://www.tng.iac.es/instruments/harps/), attached at the 3.5m Telescopio Nazionale Galileo (TNG). HARPS-N features an echelle spectrograph covering the wavelength range between 3830 to 6930 Å, with a spectral resolution of R=115,000 (sampling 3.3 px).
  • The Echelle SpectroPolarimetric Device for the Observation of Stars (ESPaDOnS888 https://www.cfht.hawaii.edu/Instruments/Spectroscopy/Espadons/) attached at the 3.6m Canada-France-Hawaii Telescope (CFHT). ESPaDOnS provides a spectral resolution of R=81,000 (sampling 0.69 px) in the wavelength range between 3700 and 10500 Å.
Typically, we obtained 1-2 epoch spectra with UVES, 2-3 with CFHT, and three or more with HARP-N. The signal-to-noise ratio (S/N) was larger than 40 for about 80% of the spectra. In total, in this work, we analysed 331 spectra for 180 DCEPs. A complete list of the time of acquisition and the individual S/N values is reported in Table 2.
Reduction of all the spectra, which included bias subtraction, spectrum extraction, flat-fielding, and wavelength calibration, was done by automatic pipelines provided by the three instrument teams so that we downloaded the science-ready one-dimensional spectra (for more details on the data reduction of HARPS-N, UVES and ESPaDOnS spectra, see Ripepi et al., 2021; Trentin et al., 2023; Bhardwaj et al., 2023, respectively).
Finally, from Table 1 it can be noted that the stars OGLE-GD-CEP-0026, OGLE-GD-CEP-1278, and OGLE-GD-CEP-1290 have been observed with both the UVES and HARP-N instruments. These stars are therefore useful for cross-checking the results obtained with the different instruments (see Sect. 4).
Figure 2: Example of spectral synthesis for our targets in the region from $\lambda$  = 6110 Å to $\lambda$  = 6151 Å. The colours differ according to the particular instrumental equipment, specifically in red the ESPaDOnS spectra, in green the HARPS spectra, and in blue the UVES spectra. The spectra are ordered, from top to bottom, by decreasing metallicity. The main spectral lines have been identified at the bottom.
@@
Figure 4: Histograms of the distribution of the chemical abundances derived in this study. Gaussian fits have been superimposed (dashed red lines) with the respective mean value and FWHM reported in each panel.
Figure 5: Chemical elements in the form of [X/Fe] plotted against the iron content. The colours differ according to the particular instrumental equipment, specifically in red the UVES spectra, in green the HARPS spectra, and in blue the ESPaDOnS spectra. The targets from R21 and T23 are depicted in orange and yellow, respectively (see also Fig. 6 and 9 for the colour legend).
Figure 6: $\alpha$-elements abundance in the form of [$\alpha$/Fe] plotted against the iron content. The colour pattern, the same as in Fig. 5, is described in the upper right legend. The horizontal dotted blue lines highlight [$\alpha$/Fe] abundances at 0.4, 0.2, and 0 dex. Typical errors for the abundances are around 0.15 dex.
@@ 3.1 Stellar parameters
A tool widely adopted in the literature for estimating effective temperature is the line depth ratio (LDR) method (Kovtyukh & Gorlova, 2000; Gray & Johanson, 1991). This method offers the advantage of being sensitive to temperature variations, while remaining unaffected by abundance changes and interstellar reddening. Typically, for each spectrum in our targets, we measured approximately 32 LDRs listed in Kovtyukh & Gorlova (2000). The mean values and weighted standard deviation were used as the final temperatures and errors.
For the remaining parameters ($\xi$ and $\log g$), we employed an iterative approach. The estimation of microturbulence involved ensuring that iron abundances showed no dependence on equivalent widths (EWs); that is, that the slope of [Fe/H] against EWs is null. To achieve this occurrence, we initially measured the EWs of 145 Fe I lines using a custom Python semi-automatic routine. The line sample was extracted from the list published by Romaniello et al. (2008), and the routine minimised errors in continuum estimation on the spectral line wings. EWs were converted to abundances using the WIDTH9 code (Kurucz & Avrett, 1981), applied to the corresponding atmospheric model calculated using ATLAS9 (Kurucz, 1993). In this calculation, the influence of $\log g$ was not considered, as neutral iron lines are insensitive to it. Subsequently, surface gravities were estimated through a similar iterative procedure, imposing ionisation equilibrium between Fe I and Fe II. The adopted list of 24 Fe II lines was extracted from Romaniello et al. (2008). Errors were estimated through the propagation of errors derived from the linear fits.
3.2 Abundances
To circumvent issues arising from spectral line blending due to line-broadening, a spectral synthesis technique was applied to our spectra. Synthetic spectra were generated in three steps: i) plane-parallel local thermodynamic equilibrium (LTE) atmosphere models were computed using the ATLAS9 code (Kurucz, 1993), employing the stellar parameters; ii) stellar spectra were synthesised using SYNTHE (Kurucz & Avrett, 1981); and iii) the synthetic spectra were convoluted to account for instrumental and line-broadening. This convolution was evaluated by matching the synthetic line profiles to a selected set of observed metal lines.
For a total of 29 different chemical elements including Fe, it was possible to detect the spectral lines used for the estimation of the abundances. For all elements, we performed the following analysis: we divided the observed spectra into intervals, 25 Å or 50 Å wide, and derived the abundances in each interval by performing a $\chi^{2}$ minimisation of the differences between the observed and synthetic spectra. The minimisation algorithm was written in Python, using the amoeba routine.
@@ 3.2 Abundances
The adopted lists of spectral lines and atomic parameters were taken from Castelli & Hubrig (2004), who updated the original parameters of Kurucz (1995). When necessary, we also checked the NIST database (Ralchenko & Reader, 2019). In Fig. 2, we plot an example of some of the spectra analysed in this work. More details about the spectral lines that can be typically detected and used to estimate chemical abundances can be found in R21 and T23. The final list of LTE abundances for all the stars is reported in Table 3. All the abundances are referred to the solar value (Grevesse et al. (2011)).
Figure 7: Example of two spectra around the Sc II line at $\lambda\lambda$ $\AA$ for the stars BQ Vel (in yellow from the T23) and ATLAS-J104.8954-08.3771 (in red from the UVES sample analysed in this paper). Other spectral lines are highlighted with dashed black lines. Atmospheric parameters and metallicity differences between the two targets are reported.
Figure 8: Example of two spectra around the Zn I line at $\lambda\lambda$ $\AA$ for the stars OGLE-GD-CEP-0228 (in yellow from the T23) and OGLE-GD-CEP-0048 (in red from the UVES sample analysed in this paper). Other spectral lines are highlighted with dashed black lines. Atmospheric parameters and metallicity differences between the two targets are reported.
@@ 4.1 Abundances results
As a first step in the analysis of our results, we verified that for the three targets observed with both UVES and HARPS-N and mentioned in Sect. 2.2 we find consistent results among the estimated abundances. In this perspective, the homogeneity in the analysis method should not be influenced significantly by the use of the different instruments to observe our targets. As is shown in Fig. 3, for all three stars we find comparable chemical values between the UVES and HARPS spectra and no systematics are detected.
@@
In Fig. 4, we show histogram distributions of each derived element abundance, for a total of 29 different chemical species. Bins have been fixed to 0.15 dex, reflecting typical error values. Each distribution has been fitted with a Gaussian curve. It is worth noting that most of these Gaussian curves have FWHM ¿ 0.25 dex, which means that for these elements we have an almost homogeneous distribution over a broad range of values. Moreover, when confronting current Fig. 4 with Fig. 3 of T23, it is possible to appreciate how this new sample better covers the whole range of abundances without sharp ‘jumps’ among bins that appeared in T23 for elements (see for example the O, Ca, and Fe panels). We complemented our sample with those presented in the previous C-MetaLL papers, R21 and T23, reaching a total of 292 pulsators analysed homogeneously.
@@
In Fig. 5, we plot each element (in its [X/Fe] form) in relation to iron. Generally, we confirm all the trends already found in R21 and T23, in which we divided the chemical species into groups according to their main channel of formation within the stellar evolution models. In more detail, for light and $\alpha$ elements (from C to Ti), we observe a descending behaviour at lower metallicities until [Fe/H] $\approx$ -0.5 dex, with a flattening effect at higher values. Moreover, we note that while in T23 the region with over solar abundances at lower [Fe/H] was poorly populated with stars that appeared as outliers (yellow points), in this work we can confirm a negative trend that flattens at higher iron abundances. In Fig. 6, we have plotted the $\alpha$ elements versus [Fe/H] (to estimate $\alpha$, we used the average between [Mg/Fe], [Si/Fe], [S/Fe], [Ca/Fe], and [Ti/Fe]), confirming the $\alpha$ enhancement at lower metallicities that had already been observed in the literature from both an observational (Hayden et al., 2015; Duffau et al., 2017; Trentin et al., 2023) and a theoretical (Tinsley, 1979; Matteucci & Brocato, 1990; Palicio et al., 2023) context. It is important to highlight that [Fe/H] and [$\alpha$/Fe] depends on both age and birth radius (see for example Wielen et al., 1996; Ness et al., 2019). Since all the objects are Cepheids, the relation illustrates the [Fe/H] dependence of [$\alpha$/Fe] at a fixed age. In other words, [Fe/H] represents in this case the (birth) Galactocentric radius rather than age.
We note that for the Sc element, most of the low-metallicity stars studied in this paper appear systematically less abundant (around solar values) and with a flatter distribution than those presented in T23, which are clearly over-abundant and with a slight descending trend towards higher [Fe/H] values. In order to investigate this discrepancy between the two samples, we selected stars with similar atmospheric parameters (and possibly similar [Fe/H] abundances) and compared the spectra around the available scandium lines. An example with the two targets, BQ Vel from T23 and ATLAS-J104.8954-08.3771, is shown in Fig. 7. It was indeed found that stars from T23 show more intense Sc lines (that is, higher abundances) than those presented in this work. This suggests that this effect could be real and not caused by systematics in the analysis. Possible effects due to different instruments are also excluded since most of the stars involved were observed with UVES.
@@
Iron peak elements (from V to Zn) present a descending trend over the range of metallicities, with the exceptions of Mn and Zn, which appear flatter. It is worth noting how for the latter element the objects presented in this work, observed with UVES (red points), populate the over-solar and solar abundance zone at lower metallicities, while most of the previous targets from T23 (also observed with UVES) were found to be under-solar. This is a similar but opposite behaviour to the Sc case. In the wavelength range covered with the UVES instrument, however, there is only one available neutral zinc line at $\lambda\lambda$ 6362.338 Å. Although this line is weak and could be difficult to detect when the S/N is not optimal, a difference in the line depth has indeed been observed as in the case of Sc (see Fig. 8). A similar study has been performed for the Copper, for which there are two neutral lines, at $\lambda\lambda$ 5105.5 and 5218.201 Å. In this case, the UVES targets are slightly under-abundant with respect to the other samples, but again a visual inspection confirmed this difference.
@@
For heavier elements (from Sr to Gd), we either find a descending behaviour (Y, Ce, Pr, Gd) or a sort of arch-form trend with a bump at [Fe/H] $\approx$--0.2 dex ( Ba, La, Nd, and Eu), after which we find that abundances decrease with increasing metallicities. This curved trend was already found in the literature related to open cluster (OC) studies (e.g. Molero et al., 2023) and is typical of those elements produced via the s-process. While the descending branch at higher metallicities is caused by the iron production from SNe Ia, at lower metallicities pollution from low- and intermediate-mass stars during the AGB phase cause the rising trend. Given the low number of stars, we cannot draw a firm conclusion for both Sr and Sm. As happened for Sc, Cu, and Zn, and also for the zirconium (Zr II line at $\lambda\lambda$ 6114.853Å), there seems to be a sort of bifurcation at the low-metallicity tail. We point out that we found the arch-form trend for some of the r-process elements as well (i.e. La and Eu).
@@
Figure 9: Galactic iron radial gradient. The colours are the same as in Fig. 5. The black line represents the linear fit computed. Results of the fit are reported both in the figure and in Table 10. Typical errors for [Fe/H] are around 0.15 dex.
Figure 10: Galactic iron azimuthal gradient for different $R_{GC}$ ranges. The colours are the same as in Fig. 5. The dashed black line represents the linear fit computed. In each panel, results of the slope are reported as well as the $R_{GC}$ range.
@@
Figure 11: Galactic radial gradient for all the estimated chemical elements. The lines and colours are the same as in Fig. 9. The linear coefficients are listed in Table 10
@@ 4.2 Radial gradients and spatial distribution
Following the same procedure described in R21 and T23 and outlined in Sect. 2.1, we estimated the Galactocentric distances, $R_{GC}$, and estimated the radial gradients for all the available chemical species. All of the fits were carried out with the python LtsFit package (Cappellari et al., 2013), which allows one to use weights on both axes and implements a robust outlier removal and error estimation of the fitted parameters. In Fig. 9, we plot the iron radial Galactic gradient. As was expected, we found a clear negative trend. Moreover, an almost homogeneous distribution in metallicity is reflected in a homogeneous distribution in terms of distance, with the stars spanning a Galactocentric distance between 5 and 20 kpc. Since in T23 we lacked a fair amount of stars with distances between 12 and 16 kpc and we had to complement the sample with literature stars, we might find different results between the current work and the previous one. One first difference is indeed the lack of a visible break at around 9.25 kpc. This is confirmed when trying to compute a two-line fit that overlaps almost perfectly with the current result. From a quantitative point of view, we did not note a big discrepancy in the value of the slope. In more detail, we found a slightly higher value (in an absolute sense) but still comparable with both the single-line case and the inner slope of the two-line fit. Similarly, when considering the results in T23 obtained after a binning division of the sample, to compensate for the non-uniform distribution, the overall slope is in agreement with our most recent study (there is a good agreement with the two-line fit as well). This puts a further emphasis on the importance of having access to as uniform and homogeneous a sample of stars as possible in terms of distance (and metallicity), one of the key points of the C-MetaLL project. Similarly to what happened in T23, we found some outliers quite far below the linear fit. These objects were observed mostly with UVES (red and yellow points) but stars observed with HARPS (orange and green points) are found as well. Since all the spectra were analysed with the same method and the measurement errors do not justify their behavior, we can consider these outliers to be real. In Fig. 11, we show the radiant galactic gradient for the other 28 elements. The estimated coefficients, together with the root mean square (rms), are listed in Table  10. These results generally agree with those listed in T23. As was already seen in T23, Ba has the largest dispersion. Other elements (i.e. Sc, Ti, and Zn) present instead a difference of $>$0.01 dex kpc-1. In these cases, we tried to repeat the linear regression including the literature sample used in T23 as well, obtaining results in perfect agreement with those of this work, once again highlighting the role of the farthest (and most metal-poor) objects and their weight in the linear fit. This is further highlighted when considering the significant difference in the number of objects used in the two works, counterbalanced by the homogeneity of the analysis and the sampling of the stars along the Galactocentric radii.
To achieve more information about the distribution in the Galaxy of our sample, we have plotted our targets in polar coordinates in Fig. 12. We computed the azimuthal gradient at different $R_{GC}$ but we did not find significant results (see Fig. 10).
@@
Figure 12: Galactic polar distribution of the stars. The new targets presented in this work are highlighted with black borders. The points are colour-coded according to [Fe/H] values and specified on the colour bar. The position of the sun is shown with a yellow-black symbol. Spiral arms from Reid et al. (2019) are superimposed. The colours are explained in the upper right legend.
Figure 13: Galactic polar distribution of the stars. The new targets presented in this work are highlighted with black borders. The points are colour-coded according to the chemical element specified on the colour bar. The position of the sun is shown with a yellow-black symbol. We alert the reader to the different colour bar scales for each element.